# CriticEval: Evaluate Large Language Model as Critic

### v4

@@ Abstract
Critique ability, i.e., the capability of Large Language Models (LLMs) to identify and rectify flaws in responses, is crucial for their applications in self-improvement and scalable oversight. While numerous studies have been proposed to evaluate critique ability of LLMs, their comprehensiveness and reliability are still limited. To overcome this problem, we introduce CriticEval, a novel benchmark designed to comprehensively and reliably evaluate critique ability of LLMs. Specifically, to ensure the comprehensiveness, CriticEval evaluates critique ability from four dimensions across nine diverse task scenarios. It evaluates both scalar-valued and textual critiques, targeting responses of varying quality. To ensure the reliability, a large number of critiques are annotated to serve as references, enabling GPT-4 to evaluate textual critiques reliably. Extensive evaluations of open-source and closed-source LLMs first validate the reliability of evaluation in CriticEval. Then, experimental results demonstrate the promising potential of open-source LLMs, the effectiveness of critique datasets and several intriguing relationships between the critique ability and some critical factors, including task types, response qualities and critique dimensions. Datasets and evaluation toolkit for CriticEval will be publicly released.
@@ 4 CriticEval Construction
Given the challenge of crafting scalar-valued and textual critiques from scratch, we construct CriticEval using a human-in-the-loop data construction pipeline as shown in Figure 2.
@@ 4.3 Reference Critique Generation and Annotation
Our empirical finding suggests that pairs of responses pose greater challenges for comparison if they perform similarly. Therefore, two kinds of pairs are first created: $(R_{\rm low},R_{\rm high})$ and $(R_{\rm med},R_{\rm high})$, designated as the easy and hard samples, respectively. Then, GPT-4-turbo is prompted to provide scalar-valued and text critiques on these pairs (Step 3 (f) in Figure 2). These outputs, labeled as “preferred label” and “text” in Figure 2, are then refined by human annotators.
@@
During human annotation, multiple human experts are asked to follow a rigorous annotation protocol, detailed in Appendix G.1, and the statistics of human annotation for reference critiques are described in Appendix G.5. Besides, several case studies are shown in Appendix H to facilitate a clear understanding of our proposed CriticEval.
@@ 5.2 Subjective Evaluation
The subjective evaluation aims to evaluate the quality of the generated textual critiques. Since responses in math reasoning and coding tasks can be verified, we only conduct subjective evaluations on other $5$ tasks for the correction dimension. In our work, GPT-4 evaluates the generated critiques by generating the chain of thought followed by the score, with our human-annotated critiques as references. It is well-known that LLMs prefer longer generations during their automatic evaluation (wang2023how, ; zeng2024evaluating, ). However, Figure 10 in Appendix I proves that there is no clue that GPT-4, with our concise and precise reference critiques as input, would give higher scores to longer critiques. The subjective scores range from $1$ to $10$. Following previous work (liu2023alignbench, ), the human-annotated reference critiques are anchored to $8$, serving as a relative scoring pivot.
@@ 5.3 Overall Score
The overall score of subjective and objective evaluation is computed as averaging on all the critique dimensions, and more details about computing the overall score can be found in Appendix J.
@@ 6 Evaluation and Analysis
The critique abilities of representative LLMs are analyzed in this section, and the overview results are shown in Table 2. Firstly, the reliability of evaluation in CriticEval are proven in Section 6.2 and Section 6.3. Then, overall analysis is described in Section 6.4. Several intriguing phenomena about some critical factors are described, including task types (Section 6.5), the response quality (Section 6.6) and the critique dimensions (Section 6.7). The complete experimental results of all evaluated LLMs on the test/dev set for each task and each critique dimension are placed in Appendix K.
@@ 6.1 LLMs to be Evaluated
35 widely used open-source and closed-source LLMs of different sizes are evaluated on CriticEval, including the instruction-tuned LLMs (qwen, ; deepseek-llm, ; cai2024internlm2, ; openai2023gpt4, ), critique-tuned LLMs that fine-tuned on critique datasets generated by GPT-4 (li2024generative, ; cui2023ultrafeedback, ; jiang2023tigerscore, ), and reward models (cui2023ultrafeedback, ; fengshenbang, ; pmlr-v162-ethayarajh22a, ). Please refer to Appendix L for all evaluated LLMs and the inference details. The prompt templates for LLMs on critique dimensions are shown in Appendix H with score rubrics listed in Figure 17 in Appendix G.2.
@@ 6.2 Reliability of Subjective Evaluation in CriticEval
Moreover, except for the feedback critique dimension, we also test the reliability of subjective evaluation on the correction and comparison critique dimensions. Specifically, we ask three human annotators to annotate the quality score of 450 critiques generated by five representative LLMs (GPT-3.5-turbo, Qwen-72B-Chat, InternLM2-20B-Chat, Mistral-7B and ChatGLM3-6B) from 9 tasks in CriticEval, and all the human annotators are guided by the same evaluation protocol in our subjective evaluation.
@@
The results are shown in Table 4. It can be found that GPT-4-turbo, with our human-annotated critiques as references, could achieve a very strong correlation with human judgments, close to the average human level. This observation proves the robust and reliable subjective evaluation of the textual critiques in the correction and comparison dimensions. Besides, the correlation scores on the correction and comparison critique dimension are higher than the feedback dimension. This phenomenon suggests that the feedback of the feedback is more challenging than the feedback of correction and comparison.
@@ 6.3 More Effective Critiques Consistently Lead to Superior Corrections
Although the reliability of subjective evaluation has been proven in Section 6.2, it is still unknown whether real feedback critique ability of LLMs is consistent with the evaluation results in CriticEval. This section answers this question by demonstrating that better critiques consistently lead to superior corrections. Specifically, two distinct LLMs are investigated: InternLM2-Chat-20B and Llama-2-70B-Chat, which exhibit notable differences in critique abilities in Table 2.
Table 5: The quality of corrections $F_{s}$ increase from top to bottom.
Models Objective Subjective
@@
InternLM2+OF 2.24 7.15 5.63 5.71
InternLM2+SF 7.53 10.33 6.85 5.80
InternLM2+HF 8.00 50.50 8.00 7.48
Llama2-70B+SF 2.24 5.33 5.63 5.54
Llama2-70B+OF 5.01 12.47 6.85 6.32
Llama2-70B+HF 8.00 42.34 8.00 7.11
Both models are prompted with three types of feedback $F_{s}$ of varying qualities: (1) Other LLM Feedback (OF), (2) LLM’s Self-generated Feedback (SF), and (3) Human-annotated Feedback (HF) which is anchored to the score of $8$, serving as a relative scoring pivot. The OF for InternLM2-Chat models originates from the Llama-2-70B-Chat, whereas for Llama-2-70B-Chat, OF is sourced from InternLM2-20B-Chat. Therefore, as shown in Table 5, for the InternLM2-20B-Chat model, the critique quality order from high to low is HF$>$SF$>$OF, which reverses to HF$>$OF$>$SF for Llama-2-70B-Chat. Note that quality scores $F_{s}$ of these critiques are evaluated within CriticEval. Subsequently, we prompt these LLMs to correct responses using these three feedback and compute both objective and subjective evaluation scores ($\boldsymbol{CR}$) within CriticEval. As illustrated in Table 5, a clear and consistent trend emerges: better feedback leads to superior corrections. This finding underscores that real critique ability of LLMs aligns closely with the evaluation results in our proposed CriticEval, i.e., critiques of LLMs with higher scores are more accurate and effective for corrections.
@@ 6.4 Overall Analysis of LLMs
As shown in Table 2, GPT-4 significantly outperforms other LLMs on most critique dimensions, while slightly underperforms our human-annotated critiques ($7.81<8$). Surprisingly, open-source LLMs are approaching state-of-the-art closed-source LLMs. For example, InternLM2-20B-Chat surpasses GPT-3.5-turbo on overall subjective scores ($6.20>5.89$). Furthermore, there is a clear relationship that the critique ability of LLMs improves steadily as the number of parameters increases (Table 2), suggesting that the critique ability of LLMs highly correlates with their capability. We also provide a clear diagram to show this relationship in Figure 12 in Appendix P. Beyond the average scores, we also categorize the textual critiques of LLMs into multiple quality intervals for more interpretable analysis, which are described in Appendix M.
@@ 6.6 Relationship with Response Quality
Before analyzing the relationship between response qualities and critique ability, it is essential to categorize the error patterns in responses. We highlight that the error patterns are related to the task type, complicating the classification of errors. To conduct a representative analysis of errors in all tasks, human annotators are asked to categorize errors into three patterns, which collectively encompass nearly all the cases: (1) Obvious error is easy to critique and correct, like apparent misuses of words in translation task; (2) Complex error is challenging to correct, regardless of whether critiques are easy to critique, like logical reasoning error in reasoning tasks; (3) Subtle error is hard to critique, while it is usually easier to revise than complex error, like slight misunderstandings of context in general chat. The distribution presented in Table 10 reveals distinct primary errors across different response qualities. More details about these error patterns in each task are described in Appendix N.
@@
Given the distribution of error patterns, we analyze critique ability of LLMs on responses with varying qualities. As shown in Table 11, high-quality responses are the hardest for feedback since they contain lots of subtle errors (Table 10). Note that the medium-quality responses have higher objective feedback scores than low-quality ones, which is inconsistent with our expectations. This phenomenon is because low-quality responses often receive very low human-annotated quality scores (near 1), while the scoring of LLMs tends to be higher, leading to a discrepancy.
For the correction dimension, low-, and high-quality responses are easier to correct than medium-quality due to the most obvious and subtle errors. There are two kinds of qualities for comparison dimension: easy and hard. Most LLMs perform better on easy samples than on hard samples. Specifically, the subjective and objective scores of easy samples are 4.78 and 39.73, respectively, higher than those of hard samples (4.55 and 29.80). For the meta-feedback dimension, LLMs achieve the highest consistency with human judgments on high-quality responses while performing worst on medium-quality responses.
@@ 7 Conclusion and Future Work
In this paper, we introduce a comprehensive and reliable benchmark for evaluating the critique abilities of LLMs, named CriticEval. Extensive evaluation first proves the reliability of CriticEval. Then, experimental results reveal the promising potential of open-source LLMs, the effectiveness of critique datasets and intriguing relationships between critique capabilities and crucial factors: task types, response qualities and critique dimensions. These observations significantly promote an in-depth understanding of the critical ability of LLMs and LLM’s self-improvement.
In the future, we plan to enhance our benchmark in several key areas: (1) Broadening the scope to include more tasks, such as tool-using; (2) Extending the benchmark to encompass other languages, like Chinese; (3) Improving the subjective evaluation protocol to allow for more fine-grained analysis; (4) Continue to evaluate LLMs and track their critique ability, like Llama-3 models; (5) Improving the quality of reference critiques by incorporating additional high-quality critiques from advanced LLMs if and only if their quality surpasses the existing reference critiques.
@@ A.1 Sub-optimal Reference Critiques
Following previous work liu2023alignbench , our work construct CriticEval with the human-in-the-loop annotation pipeline, i.e., multiple human annotators are asked to review and revise the critiques generated by GPT-4 model. Even though we have established a rigorous annotation process to ensure the high quality of annotated critiques, human annotators are inevitably influenced by GPT-4’s initial generated critiques in some more open-ended tasks, like general chat and QA tasks. This problem may result in the quality of revised critiques by human annotation still being sub-optimal. To address this problem, we plan to enhance CriticEval in our next version. Specifically, we will replace existing reference critiques with potential better critiques generated by advanced evaluated LLMs, like Claude-$3$https://www.anthropic.com/news/claude-3-family, if and only if their qualities surpasses reference critiques.
@@ A.4 More LLMs to be Evaluated
Some newly proposed LLMs and reward models lambert2024rewardbench have not been added yet, like Llama-3444https://ai.meta.com/blog/meta-llama-3/ and Claude-3 series models. Since our conclusions are summarized from evaluation results of over 35 open-source and closed-source LLMs, lacking these LLMs does not affect conclusions. We will continue to evaluate their critique ability in our future work.

### v5

@@ Abstract
Critique ability, i.e., the capability of Large Language Models (LLMs) to identify and rectify flaws in responses, is crucial for their applications in self-improvement and scalable oversight. While numerous studies have been proposed to evaluate critique ability of LLMs, their comprehensiveness and reliability are still limited. To overcome this problem, we introduce CriticEval, a novel benchmark designed to comprehensively and reliably evaluate critique ability of LLMs. Specifically, to ensure the comprehensiveness, CriticEval evaluates critique ability from four dimensions across nine diverse task scenarios. It evaluates both scalar-valued and textual critiques, targeting responses of varying quality. To ensure the reliability, a large number of critiques are annotated to serve as references, enabling GPT-4 to evaluate textual critiques reliably. Extensive evaluations of open-source and closed-source LLMs first validate the reliability of evaluation in CriticEval. Then, experimental results demonstrate the promising potential of open-source LLMs, the effectiveness of critique datasets and several intriguing relationships between the critique ability and some critical factors, including task types, response qualities and critique dimensions.
@@ 4 CriticEval Construction
Given the challenge of crafting scalar-valued and textual critiques from scratch, we construct CriticEval using a human-in-the-loop data construction pipeline as shown in Figure 2.222More considerations for utilizing human-in-the-loop annotation pipeline are described in Appendix G.
@@ 4.3 Reference Critique Generation and Annotation
Our empirical finding suggests that pairs of responses pose greater challenges for comparison if they perform similarly. Therefore, two kinds of pairs are first created: $(R_{\text{low}},R_{\text{ high}})$ and $(R_{\text{med}},R_{\text{high}})$, designated as the easy and hard samples, respectively. Then, GPT-4-turbo is prompted to provide scalar-valued and text critiques on these pairs (Step 3 (f) in Figure 2). These outputs, labeled as “preferred label” and “text” in Figure 2, are then refined by human annotators.
@@
During human annotation, multiple human experts are asked to follow a rigorous annotation protocol, detailed in Appendix H.1, and the statistics of human annotation for reference critiques are described in Appendix H.6. Besides, several case studies are shown in Appendix I to facilitate a clear understanding of our proposed CriticEval.
@@ 5.2 Subjective Evaluation
The subjective evaluation aims to evaluate the quality of the generated textual critiques. Since responses in math reasoning and coding tasks can be verified, we only conduct subjective evaluations on other $5$ tasks for the correction dimension. In our work, GPT-4 evaluates the generated critiques by generating the chain of thought followed by the score, with our human-annotated critiques as references. It is well-known that LLMs prefer longer generations during their automatic evaluation (wang2023how, ; zeng2024evaluating, ). However, Figure 7 in Appendix J proves that there is no clue that GPT-4, with our concise and precise reference critiques as input, would give higher scores to longer critiques. The subjective scores range from $1$ to $10$. Following previous work (liu2023alignbench, ), the human-annotated reference critiques are anchored to $8$, serving as a relative scoring pivot.
@@ 5.3 Overall Score
The overall score of subjective and objective evaluation is computed as averaging on all the critique dimensions, and more details about computing the overall score can be found in Appendix K.
@@ 6 Evaluation and Analysis
The critique abilities of representative LLMs are analyzed in this section, and the overview results are shown in Table 2. Firstly, the reliability of evaluation in CriticEval are proven in Section 6.2 and Section 6.3. Then, overall analysis is described in Section 6.4. Furthermore, several intriguing phenomena about some critical factors are described, including task types (Section 6.5), the response quality (Section 6.6) and the critique dimensions (Section 6.7). Finally, we elaborate and analyze the fine-grained error patterns of model-generated critiques in Section 6.8. The complete experimental results of all evaluated LLMs on the test/dev set for each task and each critique dimension are placed in Appendix L.
@@ 6.1 LLMs to be Evaluated
35 widely used open-source and closed-source LLMs of different sizes are evaluated on CriticEval, including the instruction-tuned LLMs (qwen, ; deepseek-llm, ; cai2024internlm2, ; openai2023gpt4, ), critique-tuned LLMs that fine-tuned on critique datasets generated by GPT-4 (li2024generative, ; cui2023ultrafeedback, ; jiang2023tigerscore, ), and reward models (cui2023ultrafeedback, ; fengshenbang, ; pmlr-v162-ethayarajh22a, ). Please refer to Appendix N for all evaluated LLMs and the inference details. The prompt templates for LLMs on critique dimensions are shown in Appendix I with score rubrics listed in Figure 18 in Appendix H.3.
@@ 6.2 Reliability of Subjective Evaluation in CriticEval
Moreover, except for the feedback critique dimension, we also test the reliability of subjective evaluation on the correction and comparison critique dimensions. Specifically, we ask three human annotators to annotate the quality score of 450 critiques generated by five representative LLMs (GPT-3.5-turbo, Qwen-72B-Chat, InternLM2-20B-Chat, Mistral-7B and ChatGLM3-6B) from 9 tasks in CriticEval, and all the human annotators are guided by the same evaluation protocol in our subjective evaluation. The results are shown in Table 4. It can be found that GPT-4-turbo, with our human-annotated critiques as references, could achieve a very strong correlation with human judgments, close to the average human level. This observation proves the robust and reliable subjective evaluation of the textual critiques in the correction and comparison dimensions. Besides, the correlation scores on the correction and comparison critique dimension are higher than the feedback dimension. This phenomenon suggests that the feedback of the feedback is more challenging than the feedback of correction and comparison.
@@ 6.3 More Effective Critiques Consistently Lead to Superior Corrections
Table 5: The quality of corrections $CR$ increases as the quality of feedback increases.
Models Source of Feedbacks Objective Subjective
@@
InternLM2-20B-Chat Llama2-70B-Chat 2.24 7.15 5.63 5.71
InternLM2-20B-Chat InternLM2-20B-Chat 7.53 10.33 6.85 5.80
InternLM2-20B-Chat Human-Annotated 8.00 50.50 8.00 7.48
Llama2-70B-Chat Llama2-70B-Chat 2.24 5.33 5.63 5.54
Llama2-70B-Chat InternLM2-20B-Chat 7.53 12.47 6.85 6.32
LLama2-70B-Chat Human-Annotated 8.00 42.34 8.00 7.11
Although the reliability of subjective evaluation has been proven in Section 6.2, it is still unknown whether real feedback critique ability of LLMs is consistent with the evaluation results in CriticEval. To explore this, we prompt the InternLM2-20B-Chat and Llama2-70B-Chat models to revise responses from CriticEval using three sources of feedback with varying quality levels. As illustrated in Table 5, a clear and consistent trend emerges: as the quality of the feedback increases, both the objective and subjective revision performance improves. This finding underscores that real critique ability of LLMs aligns closely with the evaluation results in our proposed CriticEval, i.e., critiques of LLMs with higher scores are more accurate and effective for corrections.
@@ 6.4 Overall Analysis of LLMs
As shown in Table 2, GPT-4 significantly outperforms other LLMs on most critique dimensions, while slightly underperforms our human-annotated critiques ($7.81<8$). Surprisingly, open-source LLMs are approaching state-of-the-art closed-source LLMs. For example, InternLM2-20B-Chat surpasses GPT-3.5-turbo on overall subjective scores ($6.20>5.89$). Furthermore, there is a clear relationship that the critique ability of LLMs improves steadily as the number of parameters increases (Table 2), suggesting that the critique ability of LLMs highly correlates with their capability. We also provide a clear diagram to show this relationship in Figure 14 in Appendix R. Beyond the average scores, we also categorize the textual critiques of LLMs into multiple quality intervals for more interpretable analysis, which are described in Appendix O.
@@ 6.6 Relationship with Response Quality
Before analyzing the relationship between response qualities and critique ability, it is essential to categorize the error patterns in responses. We highlight that the error patterns are related to the task type, complicating the classification of errors. To conduct a representative analysis of errors in all tasks, human annotators are asked to categorize errors into three patterns, which collectively encompass nearly all the cases: (1) Obvious error is easy to critique and correct, like apparent misuses of words in translation task; (2) Complex error is challenging to correct, regardless of whether critiques are easy to critique, like logical reasoning error in reasoning tasks; (3) Subtle error is hard to critique, while it is usually easier to revise than complex error, like slight misunderstandings of context in general chat. The distribution presented in Table 10 reveals distinct primary errors across different response qualities. More details about these error patterns in each task are described in Appendix P.
@@
Given the distribution of error patterns, we analyze critique ability of LLMs on responses with varying qualities. As shown in Table 11, high-quality responses are the hardest for feedback since they contain lots of subtle errors (Table 10). Note that the medium-quality responses have higher objective feedback scores than low-quality ones, which is inconsistent with our expectations. This phenomenon is because low-quality responses often receive very low human-annotated quality scores (near 1), while the scoring of LLMs tends to be higher, leading to a discrepancy. For the correction dimension, low-, and high-quality responses are easier to correct than medium-quality due to the most obvious and subtle errors. There are two kinds of qualities for comparison dimension: easy and hard. Most LLMs perform better on easy samples than on hard samples. Specifically, the subjective and objective scores of easy samples are 4.78 and 39.73, respectively, higher than those of hard samples (4.55 and 29.80). For the meta-feedback dimension, LLMs achieve the highest consistency with human judgments on high-quality responses while performing worst on medium-quality responses.
@@ 6.7 Relationship with Critique Dimensions
6.8 Fine-grained Failure Modes in Model-Generated Critiques
This section analyzes the fine-grained failure modes in model-generated critiques across feedback, comparison and correction dimensions. As illustrated in Table 13, human annotators summarize the 12 main failure modes in model-generated critiques. Then, we compute the distribution of these failure modes of all evaluated LLMs. Figure 4 demonstrate that the most frequent failure modes are missing errors (E1, E2), lacing effective comparison analysis (E7) and worse revision than references (E10) for feedback, comparison and correction dimensions, respectively. Furthermore, as shown in Figure 4, it can be observed that missing errors/suggestions (E1, E2) and inaccurate critiques (E3, E4, E8) usually lead to lower subjective scores.
Table 13: Definition of Failure Modes in Feedback, Comparison and Correction critique dimensions. E1-E6 denotes the shared failure modes of feedback and comparison dimensions, and E7-E8 belong to comparison dimension. E9-E11 belong to the correction dimension.
Critique
Dimensions
Failure Mode Description of Failure Mode
Feedback and Comparison E1 Feedback misses some errors.
E2 Feedback misses revision suggestions or suggestions are low-quality.
E3 Feedback incorrectly analyzes correct content as erroneous.
E4 Feedback content contains errors.
E5 Feedback is correct but complex.
E6
Feedback is not concise, repetitive or irrelevant.
Comparison E7 Critiques lack effective analysis between two responses.
E8 Preference between two responses is wrong.
Correction E9 Revision does not follow suggestions in feedback well.
E10
Revisions are better but have not reached the reference.
E11 There are some errors in revisions.
- Other Other Cases
Figure 3: Distribution of failure modes in each critique dimension.
Figure 4: Average subjective score of failure modes in each critique dimension.
@@ 7 Conclusion and Future Work
In this paper, we introduce a comprehensive and reliable benchmark for evaluating the critique abilities of LLMs, named CriticEval. Extensive experimental results first prove the reliability of CriticEval, and reveal the promising potential of open-source LLMs, the effectiveness of critique datasets and intriguing relationships between critique capabilities and some factors: task types, response qualities and critique dimensions. These observations significantly promote an in-depth understanding of the critical ability of LLMs and LLM’s self-improvement. In the future, we plan to enhance our benchmark in several key areas: (1) Broadening the scope to include more tasks, such as tool-using; (2) Extending the benchmark to encompass other languages, like Chinese; (3) Improving the subjective evaluation protocol to allow for more fine-grained analysis; (2) Continue to evaluate LLMs and track their critique ability, like Llama-3 models; (5) Improving the quality of reference critiques by incorporating additional high-quality critiques from advanced LLMs if and only if their quality surpasses the existing reference critiques.
Acknowledgments and Disclosure of Funding
First of all, the authors would like to thank many senior researchers for their valuable review comments provided before our submission, which greatly improved the quality of our paper: Yan Wang, Yong Hu, Rongcheng Tu, Hongli Mao, Fanshu Sun and Chen Xu. The name rankings are in no particular order. Furthermore, this work was supported by the National Natural Science Foundation of China (No. 62172039, U21B2009 and 62276110) and the MIIT Program (CEIEC-2022-ZM02-0247). The quality of corrections $CR$ increases as the quality of feedback increases.
@@ A.1 Sub-optimal Reference Critiques
Following previous work liu2023alignbench , our work construct CriticEval with the human-in-the-loop annotation pipeline, i.e., multiple human annotators are asked to review and revise the critiques generated by GPT-4 model. Even though we have established a rigorous annotation process to ensure the high quality of annotated critiques, human annotators are inevitably influenced by GPT-4’s initial generated critiques in some more open-ended tasks, like general chat and QA tasks. This problem may result in the quality of revised critiques by human annotation still being sub-optimal. To address this problem, we plan to enhance CriticEval in our next version. Specifically, we will replace existing reference critiques with potential better critiques generated by advanced evaluated LLMs, like Claude-3444https://www.anthropic.com/news/claude-3-family, if and only if their qualities surpasses reference critiques.
@@ A.4 More LLMs to be Evaluated
Some newly proposed LLMs and reward models lambert2024rewardbench have not been added yet, like Llama-3555https://ai.meta.com/blog/meta-llama-3/ and Claude-3 series models. Since our conclusions are summarized from evaluation results of over 35 open-source and closed-source LLMs, lacking these LLMs does not affect conclusions. We will continue to evaluate their critique ability in our future work.
A.5 Limited Inference Strategy
In this paper, all the evaluated models generate the critiques by using the greedy-search decoding method. There exist some inference strategies to potentially improve the model’s performance, like structured generation zheng2024sglangefficientexecutionstructured . The primary goal of our proposed CriticEval in the current stage is to construct a comprehensive and reliable benchmark for evaluating the critique ability of LLMs, and we will explore these inference strategies to improve the critique ability in our future work.